Given the information below for Seger Corporation, compute the expected share price at the end of 2017 using price ratio analysis. Assume that the historical average growth rates will remain the same for 2017. (Do not round intermediate calculations. Round your answers to 2 decimal places.)

Year 2011 2012 2013 2014 2015 2016
Price $ 91.60 $ 97.50 $ 96.20 $ 93.70 $ 115.20 $ 130.60
EPS 4.49 5.20 6.00 6.70 7.45 8.45
CFPS 7.72 8.60 8.98 10.57 12.07 13.28
SPS 47.60 52.60 52.00 55.50 66.70 74.70

Using P/E Ratio:

Using P/CF Ratio:

Using P/S Ratio:
Answer #1
Year 2011 2012 2013 2014 2015 2016 CAGR over      6 Years Ex 2017
Price $91.6 $97.5 $96.2 $93.7 $115.2 $130.6 7.35%
EPS 4.49 5.2 6 6.7 7.45 8.45 13.48% 9.59
CFPS 7.72 8.6 8.98 10.57 12.07 13.28 11.46% 14.80
SPS 47.6 52.6 52 55.5 66.7 74.7 9.43% 81.75
Ratios
P/E 20.40 18.75 16.03 13.99 15.46 15.46 -5.40% 14.62
P/CF 11.87 11.34 10.71 8.86 9.54 9.83 -3.69% 9.47
P/S 1.92 1.85 1.85 1.69 1.73 1.75 -1.90% 1.72

Compound annual growth rate(CAGR) = (((Final Value/Initial Value)^(1/number of Years) - 1) * 100)%

Therefore, expected values at the end of 2017 calculated as:- (1+CAGR of respective particular) * 2016 Values

example: (1+CAGR over 6 Years of EPS) * Price at the end of 2016

(1+0.1348)*8.45

Expected EPS at the end of 2017 = $9.59

a) Expected Price at the end of 2017 using P/E Ratio:

= Expected P/E at 2017 * Expected EPS at 2017

= 14.62 * 9.59

= $140.20

b) Expected Price at the end of 2017 using P/CF Ratio:

= Expected P/CF at 2017 * Expected CFPS at 2017

= 9.47 * 14.80

= $140.15

c) Expected Price at the end of 2017 using P/CF Ratio:

= Expected P/S at 2017 * Expected SPS at 2017

= 1.72 * 81.75

= $140.61